# Imperceptible Protection against Style Imitation from Diffusion Models

### v2

@@ Introduction
Black-box scenario. We broaden the evaluation when applied to other personalization methods. We adopt LoRA (Hu et al. 2021), a standard personalization method, and DreamStyler (Ahn et al. 2023), known for its effectiveness in style adaptation. As shown in Figure 10a, Impasto maintains robust protection effectiveness. In style protection, generalization robustness against unknown models is also a crucial aspect. To examine this, we compare by optimizing on SD v1.5 and testing on SD v2.1. As illustrated in Figure 10b, Impasto successfully preserves the robustness of PhotoGuard and AdvDM. Overall, in both facets of the black-box scenario; unknown personalization methods and diffusion models, Impasto not only maintains the generalization ability but also effectively improves protection images’ quality.
(a) Personalization methods (b) Diffusion models
Figure 10: Generalization. Impasto does not impede baselines’ generalization abilities on (a) different personalization methods; DreamBooth, LoRA, and DreamStyler or (b) model black-box scenario; SD v1.5 $\rightarrow$ v2.1.
Conclusion
We proposed Impasto to prevent style imitation in a perceptual orientation. With perceptual and difficulty-aware protections as well as perceptual constraint bank, Impasto markedly improves the quality of the protected images of the baseline protection frameworks.
Limitations. All the current protections mostly adopt adversarial perturbations, which can hamper usability due to the extensive time required for the optimization. Even accessible software (Shan et al. 2023) takes 30-60 mins to protect a $512^{2}$ image on CPU. Addressing this time constraint is a challenge that future research should aim to overcome.
Appendix
Appendix A Related Work
Image protection against generative models. The adversarial examples have also been applied to generative models to prevent the misuse of deepfake technology. For example, Ruiz, Bargal, and Sclaroff (2020a) introduced adversarial attacks against deepfake image translation systems, proposing methods such as class-transferable attacks and adversarial training to disrupt manipulated images and evade defenses. Similarly, Ruiz, Bargal, and Sclaroff (2020b) proposed a novel perturbation algorithm that effectively disrupts black-box image translation models.
More recently, the diffusion models have significantly improved their ability to generate images with high fidelity. While these developments provide substantial benefits, they have also led to more severe misuse. Consequently, recent studies on image protection have increasingly focused on the diffusion models. Most of them have explored the iterative optimization of adversarial perturbations by adapting the loss functions of diffusion models. For example, nearly all diffusion-based protection frameworks update perturbations using either the loss in Eq. 2 or Eq. 3. The former approach employs the encoder or texture loss, as seen in methods like PhotoGuard (Salman et al. 2023) and GLAZE (Shan et al. 2023). The latter adopts the diffusion or semantic loss, exemplified by frameworks such as AdvDM (Liang et al. 2023) and Anti-DreamBooth (Van Le et al. 2023). Furthermore, some approaches combine these losses (Liang and Wu 2023), while others incorporate additional losses (Shan et al. 2023; Ye et al. 2023) or logic (Van Le et al. 2023) to develop more effective protection methods.
Imperceptible adversarial examples. The concept of imperceptibility has been actively investigated in adversarial attacks (Luo et al. 2018; Guo, Frank, and Weinberger 2018; Shen et al. 2019; Sen et al. 2019; Modas, Moosavi-Dezfooli, and Frossard 2019; Croce and Hein 2019; Shamsabadi, Sanchez-Matilla, and Cavallaro 2020; Zhao, Liu, and Larson 2020; Wang et al. 2021; Luo et al. 2022; Dai et al. 2023). Among them, some studies target specific elements; e.g. the low-frequency components (Guo, Frank, and Weinberger 2018; Luo et al. 2022). Another approach use advanced constraints (Sen et al. 2019); color distance (Zhao, Liu, and Larson 2020) or quality assessment (Wang et al. 2021) are adopted. Several methods focus on restricting perturbation regions; leveraging $L_{0}$ norm to produce sparse perturbation (Modas, Moosavi-Dezfooli, and Frossard 2019; Croce and Hein 2019) or limiting perturbations to tiny salient regions (Dai et al. 2023). Our research draws inspiration from these studies. However, attacking discriminative models is fundamentally distinct from that of targeting generative models. Hence, we employ a specialized strategy designed specifically for disrupting style imitation.
Appendix B Method Details
In this section, we describe the details of Impasto— the computation and discussion of JND estimations, perceptual constraints, and the implementation specifics.
JND Estimations
  • Luminance adaptation (LA): Perturbations are less visible in regions of very low or high luminance and more noticeable in moderate lighting conditions (Jarsky et al. 2011). Hence, we modulate protection strength based on pixel luminance with a fixed adaptation model.
  • Contrast masking (CM): Perturbations can seamlessly blend into regions with intricate textures, while they leave distinct traces on flat surfaces. To simulate this, we utilize the luminance contrast (or change) of a region to measure the complexity of the pixel (Legge and Foley 1980; Wu et al. 2013a).
  • Contrast sensitivity function (CSF): Given the band-pass characteristics of the human visual system, we utilize a frequency-based JND model (Wei and Ngan 2009). The human eye is receptive to signals at modulated frequencies while exhibiting insensitivity to high-frequency components. Consequently, perturbations overlaid on high-frequency signals (e.g. edges) are less perceptible.
  • Standard deviation: To assess the spatial structure of an image, we calculate the standard deviation of local image blocks, inspired by SSIM (Wang et al. 2004). This measures the image’s structural complexity, which correlates with the sensitivity to subtle perturbations.
  • Entropy: The entropy of an image block is computed to quantify the amount of information or complexity within a local region (Wu et al. 2013b).
Luminance adaptation. The visibility thresholds within the human visual system vary to luminance levels (Jarsky et al. 2011). For instance, our eye is more sensitive under moderate lighting conditions, while discrimination is challenging in a very dark light. Similarly, we observed that perturbations are less noticeable in regions with extremely low or high luminance, prompting an increase in protection strength within these areas. To implement this concept, luminance adaptation is computed following Chou and Li (Chou and Li 1995) as in the below equation.
$LA(\mathbf{x})=\begin{cases}17\times(1-\sqrt{\frac{B(\mathbf{x})}{127}})+3,&\text{if }B(\mathbf{x})\leq 127\\ \frac{3}{128}\times(B(\mathbf{x})-127)+3,&\text{otherwise},\end{cases}$ (14)
Contrast masking. It is obvious that stimuli become less perceptible against patterned, non-uniform backgrounds. In our context, perturbations can seamlessly blend into regions with intricate textures, while they leave distinct traces on flat surfaces. To simulate this effect, we first compute the luminance contrast (Legge and Foley 1980) by convolving an input image with four directional filters. This process highlights the complexity of a region by contrasting it with its immediate surroundings. Subsequently, contrast masking is established by mapping the luminance contrast onto a logarithmic curve (Wu et al. 2013a).
where luminance contrast $LC(\mathbf{x})$ is obtained via four directional filters as $LC(\mathbf{x})=\frac{1}{16}\times\text{max}_{k=1,\dots,4}|\mathbf{x}*\nabla_{k}|$ following Wu et al. (Wu, Shi, and Lin 2019).
Contrast sensitivity function. The human visual system exhibits a band-pass response to spatial frequency and CSF represents a function of how our eye is sensitive to the contrast of signals at various spatial frequencies. High spatial frequencies correspond to rapid changes in image details, and contrast sensitivity is optimal at moderate spatial frequencies. Beyond a certain frequency threshold, known as the resolution limit, our eyes are unable to detect changes. Upon this concept, several studies have introduced and modified the CSF (Larson and Chandler 2010; Mitsa and Varkur 1993; Damera-Venkata et al. 2000; Mannos and Sakrison 1974); the CSF model $H(f,\theta)$ is given below.
where $\alpha=0.0192,\beta=0.114$ and $f$ is the radial spatial frequency, measured in cycles per degree of visual angle (c/deg). The variable $\theta$, ranging from $[-\pi,\pi]$, denotes the orientation. Additionally, $f_{\theta}=f/[0.15\times\text{cos}(4\theta)+0.85]$ accounts for the oblique effect (Larson and Chandler 2010).
The CSF model is now applied to an image in the frequency domain as $\mathbf{x}_{csf}=\mathcal{F}^{-1}[H(u,v)\times\mathcal{F}(\check{\mathbf{x}})]$, where $\mathcal{F}[\cdot]$ and $\mathcal{F}^{-1}[\cdot]$ represent the DFT and its inverse, respectively. Here, $H(u,v)$ is the DFT-version of $H(f,\theta)$, with $u,v$ being the DFT indices. The transformation from $H(f,\theta)$ to $H(u,v)$ is elaborated in Larson et al. (Larson and Chandler 2010). Prior to applying the CSF model, an input image $\mathbf{x}$ is converted into a perception-adjusted luminance form to reflect the non-linear relationship between digital pixel values and physical luminance (Larson and Chandler 2010). This is achieved by calibrating $\mathbf{x}$ to the settings of an sRGB display and converting it into perceived luminances, which indicates the relative lightness: $\check{\mathbf{x}}=\sqrt[3]{(0.02874\mathbf{x})^{2.2}}$.
Standard deviation. It has served as a crucial metric for quantifying the structural information of an image, both in pixel space (Wang et al. 2004) and feature space (Ding et al. 2020). Given that perturbations tend to be less noticeable in areas exhibiting high levels of change, (as discussed in contrast masking), we focus on these sudden changes, interpreting them as structural components. For this purpose, we compute the block-wise standard deviation in pixel space using a $9\times 9$ local block.
Post-processing All the JND estimations are min-max normalized and then inverted by subtracting from one. When constructing a perceptual map $\mathcal{M}$ with JND estimations, we encountered issues with some JNDs displaying extremely skewed distributions. Additionally, the distributions of JNDs often do not align with each other, posing challenges in their combination. While standardization aligns the distributions, we found that the continuous JND values provide weak signals as masks (e.g. on average, the protection strength drops to 50% compared to the original). To give a more distinct signal depending on the JND values, we discretize the scores by quantizing the JND values. This involves calculating JND quantiles, where the first quantile is set to 1.0. For each subsequent quantile, we multiply by a factor of $\beta$ to decrease the value in a discrete manner. We use $\beta=0.85$, resulting in quantized JND values across four quantiles as [1.0, 0.85, 0.7225, 0.6141]. This allows for a more nuanced adjustment of the protection intensity while accommodating the varied distributions of JND estimations.
Perceptual Constraints
Masked LPIPS. We use AlexNet (Krizhevsky, Sutskever, and Hinton 2012) as the backbone network for the LPIPS loss, adhering to the parameter settings in the official LPIPS implementation.
Masked low-pass. In this constraint, we utilize $\text{LP}(\cdot)$, a reconstruction function focusing on the low-frequency component. Inspired by Luo et al. (Luo et al. 2022), we implement a DWT-based reconstruction module. Given an input image $\mathbf{x}$, DWT decomposes it into one low-frequency component and three high-frequency components, as expressed by the following equation.
$\mathbf{x}_{ll}$ $=\mathbf{LxL}^{T},\;\;\mathbf{x}_{lh}=\mathbf{HxL}^{T},$ (17)
$\mathbf{x}_{hl}$ $=\mathbf{LxH}^{T},\;\;\mathbf{x}_{hh}=\mathbf{HxH}^{T}.$
Table 4: Component analysis. Within Impasto, we conduct an analysis by excluding the JND-based perceptual map generation, resulting in the creation of the perceptual map from random masks (denoted as w/o JND). Additionally, we evaluate the impact of omitting the perception-aware protection (denoted as w/o PAP).
Method Protected Image Quality Protection Performance
Impasto (Full model) 0.159 0.912 20.74 279.3
w/o JND 0.158 0.912 19.30 268.1
w/o PAP 0.199 0.890 17.57 281.1
@@ Appendix B Method Details
Table 5: Evaluating the impact of varying refinement interval step $P$. In this analysis, we assess the impact on performance by varying the refinement interval step $P$ for executing IWR and DAP.
@@
Implementation Details
When we implement Impasto into the existing protection methods, we adhere to their respective settings. Regarding the Impasto’s own components, we use following hyperparameters depicted in Eq. 13 as: $\lambda_{L}=5.0,\lambda_{LP}=10.0,\lambda_{C}=0.1$. For our PAP (Eq. 6), we utilize the $L_{\infty}$ norm. On the other hand, for the IWR loss (Eq. 7), we employ the $L_{2}$ norm. When we calculate IWR loss, since the distance between $\mathcal{L_{SP}}$ respect to $\mathcal{M}^{\prime}$ and $\mathcal{M}(\omega)$ (first term) is significantly smaller than the $L_{2}$ constraint (second term), we amplify the former term by a factor of $5\times 10^{7}$. For difficulty prediction in DAP, we set $t=5$ with $T=25$. We execute IWR and DAP every $P=4$ steps. On an NVIDIA A100 GPU, the overall perturbation optimization takes about 30 seconds for a single 512$\times$512 resolution image.
In addition, during our experiments, we observed that the magnitudes of $\mathcal{L_{E}}$ and $\mathcal{L_{SD}}$ differ considerably, with the SD loss exhibiting much smaller values. Hence, when integrating Impasto into UNet-based protection methods (e.g. AdvDM (Liang et al. 2023), Anti-DreamBooth (Van Le et al. 2023), Mist (Liang and Wu 2023)), we scale down $\lambda_{L},\lambda_{LP},\lambda_{C}$ by a factor of 0.05. This adjustment is made to balance the influence of our proposed components, ensuring that the impact of Impasto is consistent and effective in enhancing image protection.
Appendix C Experimental Settings
Datasets. In our study, we chose the painting and cartoon domains, which are most significantly impacted by style imitation from diffusion models, and used them as benchmark datasets. To create the “painting” dataset, we gather artworks from the WikiArt dataset (Tan et al. 2018). This dataset includes selections from 15 artists, 10 artworks per each. The “cartoon” dataset is sourced from the NAVER WEBTOON platform; images are cropped to contain only character depictions. This comprises 15 cartoons with 10 cartoon character images each. Following previous dataset preparation protocols (Van Le et al. 2023), each image was resized to 512$\times$512. And note that as Impasto does not require a separate training process, the data was used as benchmark set.
@@ Appendix C Experimental Settings
Table 6: Quantitative comparison of protection methods with and without Impasto. Unlike Table 1, we match the protected image quality to a similar level and compare the protection performance.
Dataset Method Protected Image Quality Protection Performance
+ Impasto 0.186 0.346 0.889 4.451 (+0.145) 22.02 (+1.03) 286.9 (+9.3)
AdvDM 0.167 0.730 0.846 3.761 (+0.000) 12.45 (+0.00) 269.0 (+0.0)
+ Impasto 0.164 0.503 0.842 3.947 (+0.231) 11.60 ( - 0.85) 276.6 (+7.6)
+ Impasto 0.241 0.894 0.779 5.480 (+0.443) 12.88 (+2.69) 162.5 (+6.6)
AdvDM 0.241 0.776 0.775 4.802 (+0.000) 10.95 (+0.00) 153.5 (+0.0)
+ Impasto 0.243 0.819 0.744 4.780 ( - 0.022) 11.90 (+0.95) 155.9 (+2.4)
Original artwork No protection Partial protection Full protection
Figure 11: Comparison of restriction types. Extending Figure 2, we compare the fine-tuning results of the original artwork with no protection, partial protection, and full protection. The red inset in the original artwork highlights the area where perturbations were applied in the partial protection case, with perturbations confined to the inset.
Evaluation. To assess the quality of the protected images, we utilize three deep learning-based full-reference quality assessments: DISTS (Ding et al. 2020), PieAPP (Prashnani et al. 2018), and TOPIQ (Chen et al. 2023). These measures compare the protected images against the original artworks. DISTS is designed to be sensitive to structural changes while exhibiting explicit tolerance to texture resampling. It achieves this by evaluating both the mean and correlation of feature maps from the compared images. PieAPP measures the perceptual error of a distorted image with respect to a reference and its training dataset. TOPIQ adopts a top-down approach, leveraging high-level semantics to direct the quality assessment network’s focus towards semantically significant local distortion regions.
To evaluate protection performance, we employ NIQE (Mittal, Soundararajan, and Bovik 2012), BRISQUE (Mittal, Moorthy, and Bovik 2011), and FID (Heusel et al. 2017). NIQE and BRISQUE are blind image quality assessment methods that only use generated images for evaluation. BRISQUE operates on the premise that distortions in natural images disrupt pixel distributions. It processes an input image and then extracts features, which are finally mapped to a mean opinion score with SVM regressor (Schölkopf et al. 2000). NIQE, in contrast, does not rely on opinion scores, instead assesses the quality of an image by comparing its statistical properties with those of a clean image dataset.
Table 7: Quantitative comparison on facial datasets.

### v4

@@ IV-B Model Analysis
TABLE V: Evaluating the impact of varying refinement interval step $P$. In this analysis, we assess the impact on performance by varying the refinement interval step $P$ for executing IWR and DAP.
@@
TABLE VI: Comparison on the blur countermeasure of PhotoGuard with and without Impasto at a similar level of protected image quality.
FID ($\uparrow$) PhotoGuard + Impasto
Clean 274.4 279.3
Blur 273.5 274.6
Difficulty-aware protection (DAP). In Fig. 9, we compare the protection performance with and without DAP evaluated on the Cartoon dataset. Fig. 9a and b show Impasto combined with PhotoGuard and AdvDM, respectively. Compared to cases without DAP, the use of DAP results in similar or improved protection efficacy for most test samples. Notably, for cases with higher protection difficulty, DAP naturally selects a stronger budget, leading to enhanced protection efficacy.
However, as DAP approximates the actual difficulty, failure cases may arise. Measuring actual protection difficulty requires: (1) performing personalization on unprotected images to obtain results, (2) protecting images with Impasto (w/o DAP), and (3) comparing the results of personalization on protected and unprotected images. This process also involves an additional perturbation optimization step based on the measured difficulty, which is impractical. Instead, DAP approximates difficulty, which can result in inaccuracies. For example, the second sample in Fig. 9a and the third and first samples from the right in Fig. 9b illustrate such cases. To evaluate the gap between actual difficulty and our approximation, we test PhotoGuard + Impasto on the Cartoon dataset using the actual difficulty measurement process. The FID score for our DAP (approximation) is 157.7, while the actual difficulty measurement yields 162.5. While the approximated approach is not perfect, it remains effective for practical purposes, given the challenges of measuring real difficulty.
Varying refinement interval step. In this analysis, we explore the effect of varying the refinement interval step $P$ on instance-wise refinement (IWR) and difficulty-aware protection (DAP) to evaluate changes in protection performance (Table V). We observe that even as $P$ increases from 1 to 4 (i.e., with less frequent refinements), both image quality and protection capability remain relatively stable. This stability can be attributed to the initial use of a JND-based perceptual map $\mathcal{M}$, which requires only minor adjustments during the IWR step. However, when $P$ is increased to 10, a significant reduction in protection performance is observed, indicating that such a large interval is insufficient for optimal refinement. Therefore, we set $P=4$ as the default, balancing performance efficacy with the number of optimization steps.
Countermeasures. To analyze the robustness of Impasto, we conduct countermeasure experiments with JPEG compression ($q=40$), Gaussian blur ($3\times 3$ kernel, $\sigma=0.02$), and Gaussian noise ($\sigma=0.02$). Results indicate a performance degradation of all protection methods when these countermeasures are applied, as they tend to remove the protective perturbations (Fig. 10). Nonetheless, Impasto demonstrates comparable robustness against such countermeasures.
For the blur countermeasure, there is a slight performance degradation with Impasto, as subtle perturbations are particularly weakened to this. However, we observed that baselines (both Photoguard and AdvDM) with low-budget protection are also vulnerable to blur. The reason for this phenomenon is that the blurring effect removes most high-frequencies, making weak perturbations less effective. We note that if baseline protection models are applied with weak protection strength (to match the protected images’ perceptual quality to ours), they also exhibit degraded protection performance.
Figure 11: Degradation tendency of PhotoGuard and AdvDM compared to Impasto as perturbation strength increases. (Top-left) Reference original image. (Top-right) Images generated after fine-tuning diffusion models on inputs with increasing levels of protection strength (more perturbations). (Bottom-right) Protected images with increasing levels of protection strength. Here, $\eta$ above generated and protected images indicates the budget, i.e., the protection strength.
(a) Personalization methods (b) Diffusion models
(c) Diffusion models with transformers
Figure 12: Generalization. Generated images when various diffusion models fine-tuned with the protected images. Note: For (a) and (b), protected image in Fig. 15 (left) is used as the input for fine-tuning. Impasto does not impede baselines’ generalization abilities on (a) different personalization methods; DreamBooth, LoRA, and DreamStyler, (b) model black-box scenario for diffusion models; SD v1.5 $\rightarrow$ v2.1, and (c) model black-box scenario for DiT; SD v1.5 $\rightarrow$ DiT-XL-2.
To demonstrate this, we match the protected image quality of PhotoGuard to PhotoGuard+Impasto and compare the protection performance in both clean (no countermeasure) and blur scenarios (Table VI). In this case, which with similar imperceptibility, Impasto shows better performance under blur, implying that such degradation is an inevitable limitation of using weaker perturbations rather than a flaw in our model design. Note that the performance gap narrows under blur because Impasto applies very weak perturbations to more noticeable regions, potentially causing minor vulnerabilities. Nevertheless, in all scenarios—whether clean or under countermeasures—Impasto demonstrates an improved trade-off between protected image quality and protection performance, highlighting its practical effectiveness.
TABLE VII: Comparison on SD with LoRA weights of PhotoGuard with Impasto at a similar level of protected image quality.
Method Protection Performance
NIQE ($\uparrow$) BRISQUE ($\uparrow$) FID ($\uparrow$)
Pretrained SD 5.620 11.46 157.7
SD w/ LoRA1 5.615 11.47 156.8
SD w/ LoRA2 5.623 11.38 155.5
  • LoRA1: https://civitai.com/models/142551/fashion-manga-style-lora
  • LoRA2: https://civitai.com/models/7801/oil-painting-style-lora
(a) Results of DISTS
(b) Results of PieAPP
(c) Results of TOPIQ
(d) Results of FID
Figure 13: Hyperparameters analysis. For PhotoGuard with Impasto, the combinations of three types of loss weights were evaluated for protection performance using FID and imperceptibility performance via DISTS, PieAPP, and TOPIQ metrics. The bar with the red border represents the results of the base settings.
Degradation tendency. Our work aims to enhance invisibility while preserving protection performance compared to baselines. To minimize perceptible degradation, the perturbation signal must be both effective and minimally intrusive. Thus, unlike adversarial attacks not tailored for invisibility, Impasto does not induce severe visual distortions. However, increasing the perturbation strength results in pronounced structural changes, as shown in Figure 11. Notably, both PhotoGuard and AdvDM with Impasto exhibit stronger reproduction of their characteristic patterns when the perturbation strength is increased (see yellow box regions).
Generalization. We broaden the evaluation when applied to other personalization methods. We adopt LoRA [18], a standard personalization method, and DreamStyler [1], known for its effectiveness in style adaptation. As shown in Fig. 12a, Impasto maintains robust protection effectiveness. In style protection, generalization robustness against unknown models is also a crucial aspect. To examine this, we compare by optimizing on SD v1.5 and testing on SD v2.1. As illustrated in Fig. 12b, Impasto successfully preserves the robustness of PhotoGuard and AdvDM. Furthermore, we conduct tests on models incorporating the recently proposed diffusion transformers (DiT) [43], and we observed that PhotoGuard (encoder-based approach) with Impasto effectively reproduces the intended target image, specifically the pattern of the grid (see Fig. 12c). Overall, in all three facets of the black-box scenario; unknown personalization methods, diffusion models, and diffusion models with transformers, Impasto not only maintains the generalization ability but also effectively improves protection images’ quality.
In addition, we conducted supplementary experiments to verify the preservation of protection performance in configurations combining pretrained SD v1.5 with finetuned weights. In this experiments, we utilize LoRA models1,2 that had been appropriately finetuned on other datasets. We evaluate whether protected images optimized for the pretrained SD maintain style protection in the finetuned versions (SD with LoRA) As listed in Table VII, it can be observed that the protection performance in terms of NIQE, BRISQUE, and FID is maintained even in the LoRA-integrated configuration.
Hyperparameters analysis. We investigate the impact of the loss weights assigned to each component of the loss function. As described in Sec.III-D, we use the following hyperparameters in Eq.14: $\lambda_{L}=5.0$, $\lambda_{LP}=10.0$, and $\lambda_{C}=0.1$. We evaluate imperceptibility and protection performance across 27 combinations by varying the three parameters at three levels each (see Figure 13). In terms of imperceptibility metrics (i.e., DISTS, PieAPP, and TOPIQ), increasing $\lambda_{L}$ generally improves performance. For FID scores, we observe noticeable fluctuations, but the default settings achieve a good balance between imperceptibility and protection. Based on this experiment, we selected the optimal hyperparameters via grid search on the validation set.
Additionally, we conduct a sensitivity analysis to evaluate the effect of diffusion process parameters (i.e., $t$) on the DAP approximation used in IMPASTO. In our default design, we fix the timestep at $t=5$ out of $T=25$ to capture local perturbation effects. This partial denoising balances computational efficiency with realism by approximating how models like DreamBooth use early-stage latent representations during fine-tuning. To assess robustness, we vary the timestep $t$ across several values (e.g., $t=4,5,6$) and observe that the imperceptibility and protection performance remained consistent.
Impasto in other domain. Although Impasto is initially proposed to prevent style imitation, its applicability can be extended beyond other domains or applications. To validate this, we conduct protection on two facial datasets, CelebA-HQ [21] and VGGFace2 [3]. As demonstrated in Table VIII and Fig. 14, adopting Impasto in these natural domains can also enhance the quality of protected images without compromising the protection performance of baseline models. In addition to facial datasets, we also conducted robustness evaluations on iCartoonFace [74], which contains a wide range of cartoon-style images, and the Stanford Online Products (SOP) dataset [57], which consists of real-world object images. As listed in the Table IX, the proposed IMPASTO demonstrates acceptable performance in terms of both DISTS and FID on diverse types of cartoon and real-world object images. This broad applicability highlights Impasto’s versatility and potential as a universal tool for protecting users’ copyright and preventing serious threats of deepfakes.
TABLE VIII: Quantitative comparison on facial datasets.